OpenStudy (anonymous):

A quadrilateral is a parallelogram if both pairs of the opposite sides are equal in length. If A(-1,0) B(3,8) C(3,2) and D are consecutive vertices of a parallelogram , what are the coordinates of D? Helpp Meee :D Thanks :D

OpenStudy (anonymous):

the coordinates of d are(-1,6)

OpenStudy (anonymous):

Can I have the Complete Solution?

OpenStudy (anonymous):

yes i'm typing

OpenStudy (anonymous):

 by this we can find the mid point x=-1+3\2=1 y=0+2\2=1